# Calibrating Translation Decoding with Quality Estimation on LLMs

### v2

@@ 4.3 Training Setups
For all experiments, we train the models using LoRA (Hu et al., 2022) with a rank of 8. We set $\alpha$ to 32 and use a dropout rate of 0.05. The training process involves a batch size of 32, a gradient accumulation step of 8, and a maximum length of 512 tokens for each sequence. We train each model for 3 epochs and select checkpoints based on the best performance measured by XCOMET on the validation set, NTREX (Federmann et al., 2022), which consists of 1997 samples for each direction. For more reliable experiments, we train models with learning rates ranging from 1e-5 to 1e-4, and report the results using the best-performing learning rate for each setting. We use Adam (Kingma and Ba, 2014) as the optimizer. Unless otherwise specified (e.g., § 2), we use CometKiwi-XXL as the signal during training and report results in all other metrics such as XCOMET, COMET, and CometKiwi-XL. All experiments are conducted using H100 GPUs, with 7B models trained on a single GPU and 13B models trained on two GPUs.
@@ 5.1 Calibration Leads to Clear Quality Improvements
Table 1: Evaluation of enxx translation on WMT24 using CometKiwi-XL and XCOMET. Results are reported for all languages covered during Tower-v1 pretraining. Note that the Tower-v2 models, including Tower-70B-v2, have not been publicly released. We report their best results as published by Rei et al. (2024a). For GPT-4o and GPT-4o-mini, we use the prompts following  Hendy et al. (2023). Results in other metrics can be found in Appendix B.1. Notably, according to Kocmi et al. (2024b), improvements of $\geq 1.99$ in XCOMET or $\geq 0.94$ in COMET scores correspond to at least 90% estimated accuracy in human judgment — both of which are achieved by our method.
@@
In this section, we demonstrate the effectiveness of our calibration approach by applying it to some state-of-the-art translation systems, e.g., Tower and ALMA series. Table 1 presents the results for Tower series under an off-policy setting (see §4.2 for the training data), measured by CometKiwi-XL, the official reference-free metric for WMT24, and XCOMET, the best performing metric as evaluated by  Freitag et al. (2023). TowerInstruct, +SFT, and +Calibration in the table denote the results decoded by beam search with a beam size of 5. TowerInstruct-7B/-13B, and TowerInstruct-Mistral-7B are official implementations (Rei et al., 2024a), supervised fine-tuned on the corresponding base models using TowerBlock, a set of high-quality translation instructions constructed through careful data selection and filtering from human-annotated translation datasets, consisting of 637k instructions.
@@
We also conducted supervised fine-tuning on the TowerBase series using 2K Best-of-N samples per direction, selected from our calibration dataset (see §4.2) based on the highest CometKiwi-XXL scores. The resulting average performances are comparable to those of official Instruction models. Note that the performance degradation when fine-tuning on the full calibration set is expected, as some of the sampled bitext examples are of low quality.
@@
When applying our calibration approach, very strong improvements can be observed across all directions, metrics, and base models, for example: (1) It leads to an average improvement of +2.8 points in KIWI-XL and +2.7 points in XCOMET over TowerInstruct-Mistral-7B. Additionally, Table 3 shows gains of +3.6 points in KIWI-XXL and +1.2 points in COMET, respectively; and (2) This performance is comparable to that of the current top-performing system, that is, Tower-70B-v2 equipped with 100-time-sampling MBR/TRR [^2] TRR (Rei et al., 2024a) denotes an ensemble strategy that applies reranking based on multiple metric model to select the best candidate from 100 sampled hypotheses. The authors report TRR results when it surpass MBR. — while being approximately 200× speed up [^3] We roughly estimate the latency of the Tower-70B-v2 model to be 10 times that of the Tower-Mistral-7B model. Meanwhile, the former employs 100× sampling, while the latter uses beam search with a beam size of 5..
@@
We also compare with CPO (Xu et al., 2024), a widely used preference optimization method in translation, following its original setting by selecting the highest- and lowest-scoring sampled candidates as accepted and rejected samples, respectively — consistent and substantial improvements over CPO are observed. Additionally, Appendix B.2 reports results based on the ALMA base model, showing gains of +2.0 XCOMET and +1.4 KIWI-XXL for out-of-English translation over CPO.
@@
Note that experimental results in on-policy settings are also provided in Appendix B.3, where substantial improvements can also be consistently observed across metrics, demonstrating effectiveness in different training dynamics. Unless otherwise specified, the following analysis focuses on the off-policy setting to simplify the investigation of the properties of our calibration method.
@@ 5.2 Calibration Leads to Direct Quality Estimation
(1) As more best-of-N samples are included in supervised fine-tuning, translation performance improves progressively. Interestingly, the corresponding quality estimation abilities, measured by Spearman scores, increase from around 51.5 to 54.0 points. We attribute this to the fact that fine-tuning on Best-of-N samples encourages the model to assign higher likelihoods to better hypotheses.
(2) Examining the effects of scaling, we observe that: (1) scaling up from 7B to 13B generally improves translation performance for both the original TowerInstruct models and the fine-tuned models; (2) however, its impact on calibration — quality estimation ability — remains minimal.
(3) Our calibration method manifests very strong improvements in both translation and quality estimation. For example, when applying our method to TowerBase-13B, the resulting model surpasses some state-of-the-art systems in both translation performance and quality estimation ability, i.e., Tower-70B-v2+MBR/TRR and CometKiwi, simultaneously.
Results in other statistics are provided in Appendix C. Overall, we observe a clear, albeit sometimes non-linear, correlation between the models’ translation performance and their quality estimation ability. These results suggest, to some extent, a unified perspective: a well-performing translation system should inherently "know" what constitutes a good translation. In turn, we also suggest optimizing translation quality by improving calibration on LLMs, rather than relying solely on extreme scaling or supervised fine-tuning, as the latter approaches show relatively limited effectiveness.
@@ 5.3 Calibration Enhances Effectiveness for Maximum a Posteriori Decoding
Figure 3: The average performance (ende,es,ru,zh) measured by COMET score for both TowerInstruct-Mistral-7B and our calibrated models, when varying sampling size for Best-of-N (BoN) sampling and applying beam search with a beam size of 5.
Compared to test-time optimization, such as Best-of-N sampling, our approach performs sampling and directly calibrates the likelihoods of hypotheses with their translation qualities during training. Ideally, for a well-calibrated model, the potential of maximum a posteriori (MAP) decoding — such as beam search — should be more effectively realized.
Figure 3 compares the effectiveness of beam search (with a beam size of 5) to that of Best-of-N sampling with varying sampling sizes, for both the baseline (TowerInstruct-Mistral-7B) and our calibrated models. To ensure fairness, we adopt a cross-metric evaluation for Best-of-N sampling: candidates are reranked using KIWI-XXL, while best-scoring results are evaluated using COMET. Two key findings are clearly illustrated in Figure 3: (1) As the Best-of-N sampling size increases, translation performance improves for both models, with our model consistently outperforming the baseline across all settings. (2) Under the MAP decoding strategy (i.e., beam search), the baseline model (dashed blue line) underperforms compared to Best-of-N sampling with a sampling size as small as 10. However, after calibration, beam search achieves performance (dashed red line) comparable to that of Best-of-N sampling with a size of 100, showing much stronger realizations of MAP capabilities. Note that beam search is dozens of times more efficient than Best-of-N sampling.
Overall, compared to the state-of-the-art translation system, our calibration method yields clear performance improvements and offers greater efficiency potential for real-world deployment. Unlike computation-intensive test-time optimization methods like Best-of-N sampling, it inherently improves effectiveness by calibrating decoding to the real-world quality objective, unlocking the potential of efficient MAP decoding rules, such as beam search.
@@ 6.1 Cross-Metric Evaluation
A potential concern with directly optimizing towards translation quality is metric gaming (Casper et al., 2023; Kovacs et al., 2024). To address this, we conduct experiments with TowerInstruct-Mistral-7B, using various metrics to assess hypothesis quality during training and to examine whether improvements are consistent across metrics. To maximize the contrast between metrics, we focus on three representative ones in our setting: (1) COMET, a strong widely used reference-based metric; (2) KIWI-XXL, the strongest reference-free metric in the KIWI series; and (3) The reference-free version of XCOMET that combines sentence-level evaluation with error span detection.
Table 2 presents the average improvements on WMT24 when applying (a) supervised fine-tuning and (b) our calibration method, using these three different metrics. It shows that: (1) When fine-tuning on our Best-of-N dataset, we generally observe slight performance degradation for all settings, except when KIWI-XXL and XCOMET are used simultaneously as both the training objective and the evaluation metric. We attribute the degradation to the fact that TowerInstruct-Mistral-7B is already stronger than GPT-4o-mini, the system we queried. Meanwhile, we attribute the score improvements in KIWI-XXL to the existence of slight metric gaming. (2) When applying our calibration methods, we observe consistently strong improvements across all training objectives and evaluation metrics, highlighting cross-metric effectiveness. (3) The reference-based metric, COMET, demonstrates relatively low effectiveness during training in both settings. Conservatively, we attribute this to the fact that reference-based metrics constrain the feasible space of positive hypotheses to be similar to the reference, thus weakening the expressiveness of the quality space.
6.2 Human Study
@@
We randomly sample 200 translation outputs in the directions of enzh/de/ru for both the baseline model (TowerInstruct-Mistral-7B) and our calibrated model from the WMT24 dataset. Several bilingual speakers, native in the respective target languages, were recruited to annotate each translation on a 0–6 scale, following the criteria outlined in (Kocmi et al., 2022), see Appendix D.1. Annotation was performed using integer scores rather than continuous real values.
Figure 4 shows the overview results for enzh annotations, covering both fine-grained score distributions and an aggregated summary. It shows that our calibrated model yields significantly more near-perfect translations (137 vs. 67 for scores of 6), with the main gains over the baseline observed in improved handling of minor-to-major errors (denoted as score of 3/4/5) in grammar, wording, or tone, see Appendix D.2. Overall, the average score for the calibrated model is 5.49 out of 6, showing very strong system effectiveness. Compared to the state-of-the-art baseline, TowerInstruct-Mistral-7B, an overwhelming win rate (106 vs. 32) can be observed. Summaries of the annotation results for other directions are in Appendix D.3. We also release all detailed human annotations to the community [^4] Detailed human annotations are available at https://huggingface.co/datasets/Calibration-Translation/Calibration-translation-human-eval..
6.3 Sensitivity to Sampling Size
@@ 7 Conclusion
This paper focuses on the well-known miscalibration problem in translation, where the training and decoding objectives possibly misalign with real-world translation quality. To mitigate this issue, we introduce a simple yet highly effective method to improve the likelihood-quality calibration during training time by directly optimizing the Pearson correlation objective. Extensive experiments demonstrate several key advantages of our method:
@@
Very clear and consistent performance improvements. Even when applied to state-of-the-art systems such as Tower, our calibration method yields substantial improvements across a wide range of automatic metrics and human evaluations. For example, a calibrated Tower-7B model with beam search achieves state-of-the-art translation performance comparable to — if not exceeding — that of the much larger Tower-70B-v2 model (which employs test-time optimizations such as MBR decoding), while the calibrated Tower-7B model requires substantially less computation.
@@
A unified view of quality optimization and estimation in translation. Our approach adopts a shared objective for both quality optimization and quality estimation in translation, offering a unified perspective on these two tasks. Our empirical analysis also supports this perspective, indicating a clear correlation between calibration level and translation performance. Experimentally, the calibrated method achieves both state-of-the-art translation performance and accurate quality estimation within a single model, with the latter even surpassing widely used QE models such as CometKiwi — all without relying on human-annotated data.
Enhanced realization for maximum a posterior decoding. By calibrating the training objective with real-world quality, the effectiveness of widely used efficient MAP decoding rules, such as beam search, is improved. This paper shows that the effectiveness of beam search originally fell far short of Best-of-N sampling in translation quality for the baseline system. However, our method not only substantially improves performance, but also enables beam search to achieve comparable performance to Best-of-N sampling with 100 hypotheses using only a beam size of 5 — dramatically reducing inference cost without sacrificing quality, showing efficiency potential for real-world deployment.
Appendix A Detailed Dataset Description

### v3

@@ 4.2 Training Setups
For all experiments, we train models using LoRA (Hu et al., 2022) with rank 8, setting $\alpha$ to 32 and dropout to 0.05. Training uses a batch size of 32, gradient accumulation of 8 steps, and sequences capped at 512 tokens. We train each model for 3 epochs, selecting checkpoints based on the best validation performance measured by XCOMET on NTREX (Federmann et al., 2022), which contains 1,997 samples per direction. To ensure robust results, we experiment with learning rates ranging from 1e-5 to 1e-4, reporting the best results for all settings. Adam (Kingma and Ba, 2014) is used as optimizer. Unless otherwise specified (e.g., § 6.1), we use CometKiwi-XXL as signal during training and report results in XCOMET, COMET, and CometKiwi-XL. All experiments use H100 GPUs, with 7B models trained on one GPU and 13B models trained on two GPUs.
@@ 5.1 Calibration Leads to Clear Quality Improvements
Table 1: Evaluation of enxx translation on WMT24 using CometKiwi-XL and XCOMET. Results are reported for all languages covered during Tower-v1 pretraining. Note that the Tower-v2 models, including Tower-70B-v2, have not been publicly released. We report their best results as published by Rei et al. (2024a). For GPT-4o and GPT-4o-mini, we use the prompts following  Hendy et al. (2023). Results in other metrics can be found in Appendix F.1. Notably, according to Kocmi et al. (2024b), improvements of $\geq 1.99$ in XCOMET or $\geq 0.94$ in COMET scores correspond to at least 90% estimated accuracy in human judgment—both of which are achieved by our method.
@@
In this section, we demonstrate the effectiveness of our calibration approach by applying it to some state-of-the-art translation systems, e.g., Tower and ALMA. Table 1 presents the results for the Tower series under an off-policy setting (see §4.2 for the training data), measured by CometKiwi-XL, the official reference-free metric for WMT24, and XCOMET, the best performing metric as evaluated by  Freitag et al. (2023). Except for closed-source models, all results are decoded by beam search with a beam size of 5. TowerInstruct-7B/-13B, and TowerInstruct-Mistral-7B are official implementations (Rei et al., 2024a), supervised fine-tuned (SFT) on the corresponding base models using TowerBlock, a set of high-quality translation instructions constructed through careful data selection and filtering from human-annotated translation datasets, consisting of 637k instructions.
@@
We also conducted SFT on the TowerBase series using 2K Best-of-N samples per direction, selected from our calibration dataset (§4.2) based on the highest CometKiwi-XXL scores. The resulting performance is comparable to the official instruction models. When fine-tuning on the full calibration set, performance is expected to degrade, as some sampled bitext examples are of low quality.
@@
When applying our calibration approach, very strong improvements can be observed across all directions, metrics, and base models. First, it leads to an average improvement of +2.8 points in KIWI-XL and +2.7 points in XCOMET over TowerInstruct-Mistral-7B. Additionally, Table 3 shows gains of +3.6 points in KIWI-XXL and +1.2 points in COMET, respectively. Second, this performance is comparable to that of the current top-performing system, that is Tower-70B-v2 equipped with 100-time-sampling MBR/TRR [^2] TRR (Rei et al., 2024a) denotes an ensemble strategy that applies reranking based on multiple metric model to select the best candidate from multiple sampled hypotheses. They report TRR results when it surpasses MBR., while being approximately 200 times faster [^3] We roughly estimate the latency of the Tower-70B-v2 model to be 10 times that of the Tower-Mistral-7B model. Meanwhile, the former employs 100× sampling, while the latter uses beam search with a beam size of 5..
@@
We also compare with CPO (Xu et al., 2024), a widely-used preference optimization method for translation, following its original setting by selecting the highest- and lowest-scoring candidates as accepted and rejected samples, respectively, achieving consistent and substantial improvements over CPO. Additionally, Appendix F.2 reports results based on the ALMA base model, showing gains of +2.0 XCOMET and +1.4 KIWI-XXL over CPO for out-of-English translation.
@@
Note that experimental results for on-policy settings are also provided in Appendix F.3, where substantial improvements can also be consistently observed across metrics, demonstrating effectiveness under different training dynamics. Unless otherwise specified, the following analysis focuses on the off-policy setting to simplify the investigation of the properties of our calibration method.
@@ 5.2 Calibration Leads to Direct Quality Estimation
(1) As more Best-of-N samples are included in SFT, translation performance progressively improves. Interestingly, the quality estimation ability (Spearman scores) increases from around 51.5 to 54.0 points. We attribute this to the fact that the model assigns higher likelihoods to better hypotheses. However, these improvements are limited and not general across languages, see Appendix G.
(2) Examining the effects of scaling, we observe that: (i) scaling up from 7B to 13B generally improves translation performance for both the original TowerInstruct models and the fine-tuned models; (ii) however, its impact on calibration, i.e., quality estimation ability, remains minimal.
(3) Our calibration method manifests very strong improvements in both translation and quality estimation. For example, when applying our method to TowerBase-13B, the resulting model surpasses some state-of-the-art systems in both translation performance and quality estimation ability, i.e., Tower-70B-v2+MBR/TRR and CometKiwi, at the same time.
Results for other statistics are provided in Appendix G. Overall, we observe a clear, albeit sometimes non-linear, correlation between the models’ translation performance and their quality estimation ability. These results suggest—to some extent—a unified perspective: a well-performing translation system should inherently ‘know’ what constitutes a good translation. In turn, we also suggest optimizing translation quality by improving calibration on LLMs, rather than relying solely on extreme scaling or supervised fine-tuning, as the latter approaches show relatively limited effectiveness.
@@ 5.3 Calibration Enhances Effectiveness for Maximum a Posteriori Decoding
Compared to test-time optimization, such as Best-of-N (BoN) sampling, our approach performs sampling and directly calibrates the likelihoods of hypotheses with their translation qualities during training. Ideally, for a well-calibrated model, the potential of maximum a posteriori (MAP) decoding, such as beam search, should be realized more effectively.
Figure 3: The average performance (ende,es,ru,zh) measured by COMET score for both TowerInstruct-Mistral-7B and our calibrated model, when varying sampling size for Best-of-N (BoN) sampling and applying beam search with a beam size of 5.
Figure 3 compares the effectiveness of beam search (with a beam size of 5) to that of BoN sampling with varying sampling sizes, for both the baseline (TowerInstruct-Mistral-7B) and our calibrated models. To ensure fairness, we adopt a cross-metric evaluation for BoN sampling: Candidates are reranked using KIWI-XXL, while the best-scoring results are evaluated using COMET. Two key findings are clearly illustrated: (1) As the BoN sampling size increases, translation performance improves for both models, with our model consistently outperforming the baseline across all settings. (2) Under the MAP decoding strategy, i.e., beam search, the baseline model (dashed blue line) underperforms compared to BoN sampling with a sampling size as small as 10. However, after calibration, beam search achieves a performance (dashed red line) comparable to that of BoN sampling with a size of 100, showing much stronger realizations of MAP capabilities. Note that beam search here is dozens of times more efficient than BoN sampling.
Overall, compared to the state-of-the-art baseline translation system, our calibration method yields clear performance improvements and offers greater efficiency potential for real-world deployment. Unlike computation-intensive test-time optimization methods like Best-of-N sampling, it inherently improves effectiveness by calibrating decoding to the real-world quality objective, unlocking the potential of efficient MAP decoding rules, such as beam search.
@@ 6.1 Cross-Metric Evaluation
A potential concern with directly optimizing towards translation quality is metric gaming (Casper et al., 2023; Kovacs et al., 2024). To address this, we conduct experiments with TowerInstruct-Mistral-7B, using various metrics to assess hypothesis quality during training and to examine whether improvements are consistent across metrics. To maximize the contrast between metrics, we focus on three representative ones in our setting: (1) COMET, a strong widely used reference-based metric; (2) KIWI-XXL, the strongest reference-free metric in the KIWI series; and (3) the reference-free version of XCOMET which combines sentence-level evaluation with error span detection.
Table 2 presents the average improvements for WMT24 when applying (a) supervised fine-tuning and (b) our calibration method, using these three different metrics. It shows that: (1) When fine-tuning on our Best-of-N dataset, we generally observe slight performance degradation for all settings, except when KIWI-XXL and XCOMET are used simultaneously as both the training objective and the evaluation metric. We attribute the general degradation to the fact that TowerInstruct-Mistral-7B is already strong, and the improvements observed in KIWI-XXL and XCOMET to the existence of slight metric gaming. (2) When applying our calibration methods, we observe consistently strong improvements across all training objectives and evaluation metrics, highlighting cross-metric effectiveness. (3) The reference-based metric, COMET, demonstrates relatively low effectiveness during training in both settings. We attribute this to the fact that reference-based metrics constrain the feasible space of positive hypotheses to be similar to the reference, thereby weakening the diversity of hypotheses recognized as valid high-quality translations.
@@
6.2 Human Study
We randomly sample 200 translation outputs in the directions of enzh/de/ru for both the baseline model (TowerInstruct-Mistral-7B) and our calibrated model from the WMT24 dataset. Several bilingual speakers, native in the respective target languages, were recruited to annotate each translation on a 0–6 scale, following the criteria outlined in (Kocmi et al., 2022), see Appendix H.1.
Figure 4 shows the results for enzh annotations, covering both fine-grained score distributions and an aggregated summary. It shows that our calibrated model yields significantly more near-perfect translations (137 vs. 67 for scores of 6), with the main gains over the baseline observed in improved handling of minor-to-major errors (denoted as score of 3/4/5) in grammar, wording, or tone, see Appendix H.2. Overall, the average score for the calibrated model is 5.49 out of 6, showing very strong system effectiveness. Compared to the state-of-the-art baseline, TowerInstruct-Mistral-7B, an overwhelming win rate (106 vs. 32) can be observed. Summaries of the annotation results for other directions are in Appendix H.3. We also release all detailed human annotations to the community  [^4] Detailed human annotations are available at https://huggingface.co/datasets/Calibration-Translation/Calibration-translation-human-eval..
6.3 Sensitivity to Sampling Size
@@ 7 Conclusions
This paper addresses the well-known miscalibration problem in machine translation by introducing a simple yet effective training-time method that directly optimizes the Pearson correlation objective to improve likelihood-quality calibration. Extensive experiments demonstrate several key advantages:
@@
Very clear and consistent performance improvements. Our calibration method yields substantial improvements across a wide range of automatic metrics and human evaluations. For example, a calibrated Tower-7B model with beam search achieves state-of-the-art translation performance comparable to—if not exceeding—that of the much larger Tower-70B-v2 model equipped with test-time optimizations technology such as MBR decoding, all while maintaining high efficiency.
@@
A unified framework for quality optimization and estimation. By using a shared objective for both tasks, our method offers a unified perspective on quality optimization and quality estimation (QE) in translation. Our empirical analysis shows a strong correlation between calibration and translation quality. Experimentally, our calibrated method achieves both top-performing translation performance and accurate quality estimation within a single model, with the latter even surpassing widely used QE models such as CometKiwi—all without relying on human-annotated data.
Enhanced realization for maximum a posterior decoding. Experimentally, we show that calibration clearly improves the effectiveness of efficient MAP decoding methods like beam search. Originally seen as less effective than Best-of-N sampling, beam search—when properly calibrated—matches the performance of Best-of-100 sampling with a beam size of just 5, significantly cutting inference cost without sacrificing quality and showing strong promise for real-world use.
Appendix A Limitations
In this paper, we examined the impact of our calibration method under an approximate maximum a posteriori (MAP) decoding (i.e., beam search) across various settings. However, we did not explore decoding with the exact highest-probability output, which is computationally intractable due to the exponential search space. We also did not investigate substantially larger beam sizes (e.g., 200), which are technically feasible but incur prohibitive computational costs. We leave these directions for future work.
Appendix B Broader Impacts
Due to resource constraints, we focus on calibrating open-source pretrained LLMs, whose language coverage is limited to that of the base models. The impact on languages beyond this scope is left for future work. Moreover, our method inherits the limitations of the underlying models. In particular, translation quality may not be consistent across languages or demographic groups, potentially raising fairness concerns. This includes the risk of amplifying societal biases, such as gender or racial bias, that may be present in the training data.
Appendix C Supplementary Details on the Methodology
C.1 Pearson Correlation Coefficient
The Pearson correlation coefficient between two sets of values $\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}$ is defined as:
$\text{Pearson}(\mathbf{x},\mathbf{y})=\frac{\sum_{i=1}^{n}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}}\sqrt{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}},$ (4)
where $\bar{x}$ and $\bar{y}$ denote the means of $\mathbf{x}$ and $\mathbf{y}$, respectively. Let us define the mean-centered and $\ell_{2}$-normalized versions of $\mathbf{x}$ and $\mathbf{y}$ as:
$\tilde{\mathbf{x}}=\frac{\mathbf{x}-\bar{x}}{\|\mathbf{x}-\bar{x}\|},\quad\tilde{\mathbf{y}}=\frac{\mathbf{y}-\bar{y}}{\|\mathbf{y}-\bar{y}\|}.$ (5)
Then, the Pearson correlation simplifies to the dot product:
$\text{Pearson}(\mathbf{x},\mathbf{y})=\tilde{\mathbf{x}}^{\top}\tilde{\mathbf{y}}.$ (6)
This shows that the Pearson loss can be computed using only mean-centering, normalization, and a single dot product—operations that are both differentiable and computationally efficient.
Appendix D Detailed Dataset Description